what is the solution set for the system

9x-12y=1 and -3x+4y=27

Answer by oscargut(891)

9x-12y=1
-3x+4y=27 (x3)
9x-12y=1
-9x+12y=81
adding
0x+0y=81
so the system is inconsistent (has no solutions)